OpenStudy (anonymous):

partial derivative of xysin^-1(yz) with respect to y.

OpenStudy (anonymous):

ah. so you have to use chain rule too? not just product rule?

OpenStudy (anonymous):

treat all other variables as constants. using chain rule and skipping steps, simplifying early, ect, i get: \[x(\sin^{-1}(yz) + \frac{ yz }{ \sqrt{1 - (yz)^{2}} })\]

OpenStudy (anonymous):

sorry. i mean product rule, not chain rule.

OpenStudy (anonymous):

Hmm. could do you use chain rule when there are 3 things to multply? x*y*sin^-1(yz)

OpenStudy (anonymous):

meant product rule*

OpenStudy (anonymous):

you could use product rule when you have multiple things, but in this case you only have two functions of y

OpenStudy (anonymous):

how do you use the product rule is what i meant to say

OpenStudy (anonymous):

the product rule for 3 + functions gets messy

OpenStudy (anonymous):

for f(y)g(y) its f'(y)g(y) + f(y)g'(y)

OpenStudy (anonymous):

Alright, i'll look up how to do it.

OpenStudy (anonymous):

\[\frac{ d }{ dy } f(y)g(y) = f'(y)g(y) + f(y)g'(y)\] where: \[f'(y) = \frac{ d }{ dy } f(y)\]

OpenStudy (anonymous):

our case we have:\[f(y) = xy\] \[g(y) = \sin^{-1}(yz)\]

OpenStudy (anonymous):

finding f'(y) and g'(y) before jumping in and plugging is helpful for harder functions

OpenStudy (anonymous):

\[f'(y) = x\] \[g'(y) = \frac{ z }{ \sqrt{1 - (yz)^{2}} }\]

OpenStudy (anonymous):

i found a video that explained it nicely, you dont have to post anything more. thanks for clarifying

OpenStudy (anonymous):

np. glad you know it :)

OpenStudy (anonymous):

ty once again